a. How many moles of Ar are present in a 2.7 L flask that has a pressure of 1.02 atm at a temperature of 26°C?
1
Expert's answer
2018-10-16T05:27:08-0400

Answer:

T=26+273=299 K

V=2,7 L

P=1,02 atm

R=0.082

Formula: PV=nRT, n=PV/RT

n=PV/RT =1,02 x 2,7 / 0,082 x 299= 0.1123 mol.

n= 0.1123 mol (Ar)